Given f(x) = -20 − 2x, find a number, c, in the domain such that f(c) = -14. Write the corresponding ordered pair of the function.
c =
ordered pair (x, y) = ( , )

Answer by jim_thompson5910(35256)


f(x) = -20 - 2x

f(c) = -20 - 2c

-14 = -20 - 2c

-14 + 20 = -2c

6 = -2c

-2c = 6

c = 6/(-2)

c = -3

So the ordered pair is (-3, -14)